In a study of the life expectancy of 450people in a certain geographic region, the mean age of death was 75 years, and the standard deviation was 5.4 years. If a sample of 45 people from this region is selected, what is the probability that the mean life expectancy will be less than 69



years

1
Expert's answer
2022-04-25T13:42:28-0400

We have a normal distribution, "\u03bc=75,\u03c3=5.4,n=45."

Let's convert it to the standard normal distribution,

"\bar{z}=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\cfrac{69-75}{5.4/\sqrt{45}}=-7.45,\\
P(\bar{X}<69)=P(\bar{Z}<-7.45)=\\
=0\text{ (from z-table).}"